THE EXTRAORDINARY TRANSMISSION OF METALLIC ARRAYS OF SUBWAVELENGTH APERTURES FOR THE MOLECULES OF THE CELL MEMBRANE

SHANNON TEETERS-KENNEDY, KENNETH R. RODRIGUEZ, SHAUN M. WILLIAMS, ALEXANDRA SUDNITSYN, SUMMIT SHAH, FRANK HROVAT, and \underline{JAMES V. COE}, 100 W. 18th Ave., The Ohio State University, Columbus, OH 43210-1173..

Ebbesen’s extraordinary transmission effect^{a} involves passing more light through apertures in a metal film than is incident upon the holes themselves. We have moved this effect into the infrared region (by using metal arrays with hole widths of about 3 microns and hole-to-hole spacings of about 13 microns) in order to get transmission over the traditional range of molecular vibrations. A method has been developed to record enhanced infrared absorption spectra of self-assembled monolayer and bilayer coatings on metal with 100 to 1000-fold enhancements in absorbance over literature reports. The enhanced infrared absorption spectra of alkanethiolate, phospholipid, and cholesterol assemblies will be presented.

aT. W. Ebbesen, H. J. Lezec, H. F. Ghaemi, T. Thio, and P. A. Wolff, Nature (London) \underline{391}(6668), 667 (1998).